OpenStudy (anonymous):

Write the standard form of the line that passes through the given points, (-1, -3) and (2, 1)???

OpenStudy (lochana):

simplify this \[\frac{ y - 1 }{ x - 2 } = \frac{ 1+3 }{ 2+1 }\]

OpenStudy (lochana):

http://www.coolmath.com/algebra/08-lines/12-finding-equation-two-points-01 read this to find why it works that way

OpenStudy (anonymous):

Oh, that might make more sense then the lesson i was given. anyway, I don't know if I was supposed to really simply that but i got 4/3

OpenStudy (lochana):

yes. that's right. final answer will be 3y = 4x - 5

OpenStudy (anonymous):

Okay...thank you for the link as well.

Directrix (directrix):

Remember to write answer in standard form: Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers.

Directrix (directrix):

>>final answer will be 3y = 4x - 5 In standard form, 4x - 3y = 5 @actunnatural